# Bridging the Gap Between Promise and Performance for Microscaling FP4 Quantization

### v1

@@ 4.2 GPU Kernel Support for MR-GPTQ via QuTLASS
While the product $WH_{k}$} is pre-fused in the weights, $XH_{k}$} occurs online. To avoid diminishing the benefits of FP4 hardware acceleration, QuTLASS provides lightweight fused kernels for online rotation. These kernels support “unimodal” $k\times k$ block diagonal matrices with $k\in\{16,32,64,128\}$. We found experimentally that, for $k<256$, dense transformations remain memory-bound, meaning that any rotation (not just Hadamards) can be applied at essentially the same cost, as the full matrix can be loaded at runtime. (However, we found the Hadamard transform to yield best accuracy results, e.g., see Tables 12 and 13). To further reduce overhead, quantization and scale calculation are fused into the transformation kernel as a custom epilogue function. QuTLASS currently supports MSE and Abs-Max quantization methods, while its template-based design allows new methods to be easily integrated.
@@ 5.1 Experiments with Emulated Quantization
As a control, we also implement weight-only quantization, via RTN, GPTQ, AWQ [34], as well as Hadamard rotations followed by RTN, denoted as RTN + HT. These results closely follow our findings for W&A quantization, and are thus deferred to the Appendix. In Appendix G, we perform an exhaustive sweep over DCT, DST, Hadamard, and GSR transforms and block sizes showing that the Hadamard transform matching the quantization group size provides the best results on average.
@@ 6 Conclusion
The code for reproducing the quantization and accuracy evaluation experiments is provided at https://github.com/IST-DASLab/FP-Quant. The QuTLASS code is at https://github.com/IST-DASLab/qutlass.
@@ Appendix C Scale quantization analysis
MXFP format adopts E8M0 grid with exponentially spaced levels. It allows to represent values with very small and large magnitude, yet the distance between adjacent levels can be pretty large resulting in large approximation errors. E4M3 grid used in NVFP, on the other hand, has much narrower dynamic range [448,448][-448,448] with levels spread more uniformly. We note, that the sign bit is in fact never utilized, given that the scale is a non-negative value by definition.
@@ Appendix E QuTLASS results on GeForce GPUs
Appendix F Standard deviation
@@
Ident+MinMax+DefOrder 89.16 0.372 - -
@@
Appendix G The Effect of Different Linear Transforms

### v2

@@ 2 Background on Microscaling Floating-Point Formats
Figure 1: Schematic illustration of the MXFP4 (left) and NVFP4 (right) microscaling formats.
@@ 4.1 Adapting GPTQ to FP4 Formats
MXFP scale fitting
Additionally, for MXFP in Appendix F we propose a simple scale-fitting modification, mapping the excessively large E8M0 range into data range that significantly improves performance of MXFP quantized models.
@@ 4.2 GPU Kernel Support for MR-GPTQ via QuTLASS
While the product $WH_{k}$} is pre-fused in the weights, $XH_{k}$} occurs online. To avoid diminishing the benefits of FP4 hardware acceleration, QuTLASS provides lightweight fused kernels for online rotation. These kernels support “unimodal” $k\times k$ block diagonal matrices with $k\in\{16,32,64,128\}$. We found experimentally that, for $k<256$, dense transformations remain memory-bound, meaning that any rotation (not just Hadamards) can be applied at essentially the same cost, as the full matrix can be loaded at runtime. (However, we found the Hadamard transform to yield best accuracy results, e.g., see Tables 13 and 14). To further reduce overhead, quantization and scale calculation are fused into the transformation kernel as a custom epilogue function. QuTLASS currently supports MSE and Abs-Max quantization methods, while its template-based design allows new methods to be easily integrated.
@@ 5.1 Experiments with Emulated Quantization
As a control, we also implement weight-only quantization, via RTN, GPTQ, AWQ [34], as well as Hadamard rotations followed by RTN, denoted as RTN + HT. These results closely follow our findings for W&A quantization, and are thus deferred to the Appendix. In Appendix H, we perform an exhaustive sweep over DCT, DST, Hadamard, and GSR transforms and block sizes showing that the Hadamard transform matching the quantization group size provides the best results on average.
@@ 6 Conclusion
The source code required to reproduce the quantization and accuracy‑evaluation experiments is available in the FP-Quant repository. The implementation of QuTLASS can be accessed here.
Models.
MR‑GPTQ quantized models are hosted in a dedicated Hugging Face collection.
@@ Appendix C Scale quantization analysis
MXFP format adopts E8M0 grid with exponentially spaced levels. It allows to represent values with very small and large magnitude, yet the distance between adjacent levels can be pretty large resulting in large approximation errors. E4M3 grid used in NVFP, on the other hand, has much narrower dynamic range [448,448][-448,448] with levels spread more uniformly. We note that the sign bit is never utilized, since the scale is non-negative by definition. We experimented with asymmetric quantization of scales; however, this led to significantly worse results because most scales (which have small magnitudes) were mapped to the vicinity of the smallest negative value, where quantization levels are sparse. A reasonable alternative would be to map the sign bit to either the exponent or mantissa.
@@ Appendix E QuTLASS results on GeForce GPUs
Appendix F MXFP scale fitting
The principal cause of the poor performance of MXFP quantization is the large‑scale quantization error. On the one hand, the E8M0 format allows representing extremely small values, such as $2^{-127}$, and extremely large values, such as $2^{128}$. On the other, as shown in Figure 4, the actual range of weights and activations is much more narrow, making the E8M0 grid too coarse. To address this mismatch, we propose a simple modification: fit the quantization grid to the data range.
The original MXFP quantization grid, with 4/3 re-scaling, quantizes scales as follows:
$s_{\mathrm{E8M0}}=(4/3)\cdot 2^{\mathrm{clamp}(\mathrm{round}(\log_{2}s),-128,127)}.$ (1)
One can estimate the smallest $s_{\min}$ and the largest $s_{\max}$ values in given tensors, such that the smallest value is mapped to -128, and the largest to 127:
$s_{\mathrm{E8M0}}=2^{(\log_{2}s_{\max}-\log_{2}s_{\min})\mathrm{clamp}(\mathrm{round}\left(255\cdot\frac{\log_{2}s-\log_{2}s_{\min}}{\log_{2}s_{\max}-\log_{2}s_{\min}}\right),0,255)+\log_{2}s_{\min}}.$ (2)
The modification reduces the scale‑quantization error, by rescaling the exponent. In effect, we replace the exponent by a value $2^{\alpha},\alpha\in(0,1)$, followed by rescaling:
$s=2^{\alpha q+\beta}$ (3)
We provide the comparision between the original MXFP and modified version, called MXFP4, in Table 11. In almost every setting, MXFP4 yields a substantial performance boost relative to vanilla MXFP4 and performs close to NVFP4. One should note that MXFP4 requires 4.25 bit per parameter compared to 4.5 for NVFP.
Method Format Llama3 (8B) Qwen3 (8B)
RTN MXFP4 87.8 93.7
MXFP4 94.3 (+6.5) 96.3 (+2.6)
NVFP4 94.7 98.9
RTN + HT MXFP4 89.2 93.6
MXFP4 93.9 (+4.7) 96.3 (+2.7)
NVFP4 93.8 96.0
GPTQ MXFP4 89.5 94.1
MXFP4 95.2 (+5.7) 92.3 (-1.8)
NVFP4 95.7 98.1
MR-GPTQ MXFP4 93.6 95.2
MXFP4 94.9 (+1.3) 98.5 (+3.3)
NVFP4 95.8 97.4
Table 11: Per-model recoveries with vanilla MXFP4 format and the proposed modification denoted by MXFP4.
Efficient arithmetic with MXFP scales requires that the weight and activation tensors in a layer share the same exponent. Consequently, the product of the activation scale $s_{A}$ and the weight scale $s_{W}$ can be expressed as:
$s_{A}s_{W}=2^{\alpha_{A}q_{A}+\beta_{A}}2^{\alpha_{W}q_{W}+\beta_{W}}=2^{\alpha_{A}q_{A}+\beta_{A}+\alpha_{B}q_{B}+\beta_{B}}.$ (4)
In the above expression, $\alpha_{A}$ and $\alpha_{B}$ have to the same for one to express scale multiplication in terms of power addition.
Appendix G Standard deviation
@@
Ident+MinMax+DefOrder 89.16 0.372 84.85 1.95
@@
Appendix H The Effect of Different Linear Transforms